OpenStudy (anonymous):

Find the sum of the infinite geometric series, if possible. 128 + 64 + 32 + . . .

OpenStudy (anonymous):

128|+64+32+...=2^7+2^6+2^5+...

OpenStudy (anonymous):

geometric series with a=128 and q=1/2